OpenStudy (anonymous):

if I walk 10,560 feet per hour, how long will a 10,100 feet trail take if I walked it?

OpenStudy (nincompoop):

10,100 what? meters, feet, miles, kilometers? be specific

OpenStudy (nincompoop):

\[10,100ft \times \left( \frac{ 1hr }{ 10,560ft } \right)=?\]

OpenStudy (anonymous):

ok thanks

OpenStudy (nincompoop):

you're welcome

OpenStudy (nincompoop):

close this question when you're done analyzing it